OpenStudy (summersnow8):

Verify identity: csc^2 beta - cos^2 beta - sin^2 beta = cot^2 beta

hartnn (hartnn):

csc^2 beta - cos^2 beta - sin^2 beta =csc^2 beta - [cos^2 beta + sin^2 beta] got this step, i just factored out - *minus*

OpenStudy (summersnow8):

okay

hartnn (hartnn):

now you know what cos^2 beta + sin^2 beta simplifies to ?

OpenStudy (summersnow8):

1

OpenStudy (summersnow8):

so csc^2 beta -1 = cot^2 beta

OpenStudy (whpalmer4):

\[\csc^2\beta-\cos^2\beta-\sin^2\beta = \cot^2\beta\] \[\csc^2\beta -(\cos^2\beta+\sin^2beta) = \cot^2\beta\] \[\csc^2\beta-1=\cot^2\beta\]

OpenStudy (whpalmer4):

and that's probably on your identity sheet, right?

hartnn (hartnn):

yes, @Summersnow8 , correct :)

OpenStudy (summersnow8):

but i cant use the sheet

OpenStudy (summersnow8):

@hartnn then what

hartnn (hartnn):

csc^2 beta -1 = cot^2 beta means you have verified your given identity, you proved left side = right side

OpenStudy (whpalmer4):

\[\cot^2\beta = \frac{1}{\tan^2\beta} = \frac{1}{\frac{\sin^2\beta}{\cos^2\beta}} = \frac{\cos^2\beta}{\sin^2\beta}\] and \[\csc^2\beta = \frac{1}{\sin^2\beta}\]so you have \[\frac{1}{\sin^2\beta}-1=\frac{\cos^2\beta}{\sin^2\beta}\] \[\frac{1}{\sin^2\beta}-\frac{\sin^2\beta}{\sin^2\beta}=\frac{\cos^2\beta}{\sin^2\beta}\] and using an identity you already used, you can see that is true...

OpenStudy (whpalmer4):

(look at the numerators)

hartnn (hartnn):

you still have doubts ?

OpenStudy (summersnow8):

yeah, but it is okay. I can't make sense of these problems, 7 hours...... im ready to just throw in the towel

OpenStudy (whpalmer4):

1-sin^2 = cos^2 right? cos^2 + sin^2 = 1

hartnn (hartnn):

but you have already verified it completely, let me list out the steps \(left \:\:side =\csc^2\beta-\cos^2\beta-\sin^2\beta \\ =\csc^2\beta -(\cos^2\beta+\sin^2 \beta) =\csc^2 \beta-1 \\= \cot^2\beta = right \:\:side \) hence verified

OpenStudy (summersnow8):

i cant think anymore.

OpenStudy (whpalmer4):

I know the feeling :-)

OpenStudy (summersnow8):

thanks anyway

OpenStudy (whpalmer4):

take a fresh look at it tomorrow.

OpenStudy (summersnow8):

it is due tomorrow, can't

OpenStudy (whpalmer4):

okay, later tonight. take a 15 minute brisk walk. light exercise helps...

OpenStudy (summersnow8):

i rather just give up, this is stupid